Tulu is 24 years older than Hadas six years ago. Tulu was five times as old as Hadas was at the time. Find the present ages of Tulu and Hadas.
Answer by palanisamy(496)

Let x be the age of Hadas at present.
Tulu was five times as old as Hadas was at the time.
Therefore the age of Tulu at present is 5x.
Tulu is 24 years older than Hadas six years ago
5x-24-6 = x-6
4x = 24
x = 24/4
x = 6
Therefore, at present the age of Hadas =6
at present the age of Tulu =30